OpenStudy (anonymous):

Can anyone help me with only 5 questions of exponential decay? Pleaseeee i'll give medals :(

jhonyy9 (jhonyy9):

goooo

jhonyy9 (jhonyy9):

where are ???

OpenStudy (anonymous):

why is your picture a pic of your butt?

OpenStudy (anonymous):

yummmm

OpenStudy (anonymous):

A farmer in china discovers a mammal hide that contains 71% of its original amount of C-14. Find the age of the mammal hide to the nearest year.

OpenStudy (anonymous):

do you know the half life or carbon 14?

OpenStudy (bahrom7893):

e^(-kt) = 0.41

OpenStudy (bahrom7893):

wait sorry e^(-kt) = 0.71

OpenStudy (bahrom7893):

kt = ln(1/0.71) = ln(1.408)

OpenStudy (anonymous):

i think it is 5700 years. if so solve \[.71=(\frac{1}{2})^{\frac{t}{5700}}\] for t via \[\ln(.71)=\frac{t}{5700}\ln(.5)\] \[t=5700\times \frac{\ln(.71)}{\ln(.5)}\]

OpenStudy (anonymous):

i get 2816 years

OpenStudy (bahrom7893):

k= 0.0001 t = ln(1.408)/0.0001 = 3421

OpenStudy (bahrom7893):

I got that, but I think satellite must be right

OpenStudy (anonymous):

no k is not .0001 that is rounding error

OpenStudy (bahrom7893):

oh okay then satellite is right, i dunno the half life of C-14 just googled it and that was what i got

OpenStudy (anonymous):

it is easier in these problems to work with half lives.

OpenStudy (anonymous):

So it's 2816 years?

OpenStudy (anonymous):

that is what i got, yes.

OpenStudy (anonymous):

check my calculator work though